OpenStudy (anonymous):

Two investments totaling $17,500 produce an annual income of $825. One investment yields 4% a year, while the other yields 5% per year. How much is invested at each rate?

OpenStudy (anonymous):

\[x + y = 17 500\]\[0.04x + 0.05y = 825\] Solve for either x or y...

OpenStudy (anonymous):

that will work. so will \[.04x+.05(17500-x)=825\] depending on whether you like one variable or two

OpenStudy (anonymous):

So how much is invested at 4% and how much is invested at 5%?

OpenStudy (llort):

rearrange the equations to isolate a variable then replace the variable in one equation with the information from the other.